# Exponential Convergence of Augmented Primal-dual Gradient Algorithms for Partially Strongly Convex Functions

### v2

@@ II Preliminaries
The matrix $T$ is of full row rank and satisfies $\kappa_{1}I_{m}\preceq TT^{\top}\preceq\kappa_{2}I_{m}$ for some $\kappa_{1}$, $\kappa_{2}>0$.
@@ III Augmented Primal-dual algorithm
The augmented Lagrange for problem (1) is
@@
$\dot{x}=$ $-\nabla f(x)-T^{\top}\lambda-\alpha T^{\top}(Tx-b)$ (4a)
$\dot{\lambda}=$ $Tx-b.$ (4b)
@@
$\begin{bmatrix}\dot{\tilde{x}}\\ \dot{\tilde{\lambda}}\end{bmatrix}=\begin{bmatrix}-\alpha T^{\top}T&-T^{\top}\\ T&\mathbf{0}\end{bmatrix}\begin{bmatrix}\tilde{x}\\ \tilde{\lambda}\end{bmatrix}:=A\begin{bmatrix}\tilde{x}\\ \tilde{\lambda}\end{bmatrix}$ (5)
where $\tilde{x}=x-x^{*}$ and $\tilde{\lambda}=\lambda-\lambda^{*}$. It can be shown that $A$ is Hurwitz if $m=n$.
@@
This result shows that exponential convergence can be obtained when $m=n$ without any requirement on the strong convexity. This implies that we only need partially strong convexity as stated in Assumption 4 to enforce exponential convergence of the states “irrelevant” to the equality subspace when $m\leq n$.
@@ V-B Relaxing global strong convexity
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x_{p})\|^{2}$ (19)

### v3

@@ II Preliminaries
The matrix $T$ is of full row rank and satisfies $\kappa_{1}I_{m}\preceq TT^{\top}\preceq\kappa_{2}I_{m}$ for some $\kappa_{2}\geq\kappa_{1}>0$.
@@ III Augmented Primal-dual algorithm
The augmented Lagrangian for problem (1) is
@@
$\dot{x}=$ $-\nabla f(x)-T^{\top}\lambda-\alpha T^{\top}(Tx-b)$ (4)
$\dot{\lambda}=$ $Tx-b.$
@@
$\begin{bmatrix}\dot{\tilde{x}}\\ \dot{\tilde{\lambda}}\end{bmatrix}=\begin{bmatrix}-\alpha T^{\top}T&-T^{\top}\\ T&\mathbf{0}\end{bmatrix}\begin{bmatrix}\tilde{x}\\ \tilde{\lambda}\end{bmatrix}$ (5)
where $\tilde{x}=x-x^{*}$ and $\tilde{\lambda}=\lambda-\lambda^{*}$. It can be shown that the system matrix in (5) is Hurwitz if $m=n$.
Let us first introduce the following lemma that shows the Hurwitz property for matrices in a specific form.
@@
This result shows that when $m=n$, exponential convergence for system (5) can be obtained by letting $F=T^{\top}T$ in Lemma 1, without any requirement on the strong convexity. This implies that we only need partially strong convexity as stated in Assumption 4 to enforce exponential convergence of the states “irrelevant” to the equality subspace when $m\leq n$.
@@ V-B Relaxing global strong convexity
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x_{p})\|^{2}.$ (19)